# Enabling local editing in Diffusion Models

### v1

@@
A Algorithms & Implementation Details
@@ A.1 Algorithms
Computing Joint and Individual Components in the Latent Space of DMs: Algorithm 1 summarizes our method for decomposing the Jacobian of each region into a joint and an individual component. We start with a real image $\mathbf{x}_{0}$, a denoising network $\mathbf{e}_{\theta}$ and a set of regions $M=\{\mathbf{m}_{1},\cdots,\mathbf{m}_{N}\}$, along with the denoising timestep $t$ and joint and individual ranks $r_{C}$ and $r_{A}$. Starting with $\mathbf{x}_{0}$, we obtain its corresponding noise $\mathbf{x}_{T}$ via DDIM Inversion [song2020denoising]. We denoise $\mathbf{x}_{T}$ up to timestep $t$ with the standard DDIM reverse process (see Eq. (3) in text). Then we obtain the SVD of the Jacobian of each region with the Subspace Iteration (Algorithm 2). Finally, by utilizing the JIVE (Algorithm 3) we obtain the joint and individual components $\mathbf{C}$ and $\mathbf{A}_{i}$ for each region $i$. The rows of $\mathbf{A}_{i}$ are latent directions that result in meaningful local edits within region $i$.
@@
2:     $\mathbf{x}_{T}\leftarrow\text{DDIMInversion}(\mathbf{x}_{0})$triangleright DDIM Inversion [song2020denoising]
@@
Jacobian Subspace Iteration: As discussed in Section 4.2 the dimension of the latent space and the output image, result in a Jacobian matrix of approximately 6B parameters. To efficiently calculate the SVD of the Jacobian of the denoising network without storing it in memory we rely on Jacobian Subspace Iteration as proposed by Haas et al. [haas2023discovering]. Note that differently from [haas2023discovering] we calculate the SVD of the Jacobian of the denoising network explicitly within a region of interest $\mathbf{m}$.
@@
Joint and Individual Variation Explained (JIVE) Algorithm: For completeness we provide the iterative JIVE algorithm [lock2013joint]. Given a set of $N$ matrices $\mathbf{X}=[\mathbf{X}^{(1)}\cdots\mathbf{X}^{(N)}]$, JIVE iteratively approximates their joint and individual components.
@@ A.2 Impementation Details
Asyrp For Asyrp [kwon2022diffusion] we use the author’s official code [^1] https://github.com/kwonminki/Asyrp_official. Since Asyrp is a supervised approach we train each attribute with 1k samples following the recipe detailed in their paper. Asyrp uses CLIP to obtain the editing directions and requires a source caption $y^{\text{source}}$ and a target caption $y^{\text{target}}$. The source and target captions we use to obtain the editing directions for each attribute are:
@@
Haas et al. We implement the method of Haas et al. [haas2023discovering] following their paper  [^2] At the time of writing the paper the official codebase of Haas et al. was not public yet.. The editing directions for Smile, Close Eyes, Red Lips, and Gaze edits were obtained from CelebA-HQ samples with index numbers 00009, 00000, 00042, 00003 respectively.
@@
B Additional Experimental Resuls
@@ B.4 Additional Comparisons
In Figure 5 we present additional qualitative comparisons for two semantic directions not shown in the main paper. Asyrp [kwon2022diffusion] produces considerable distortions and artifacts in the edited regions and in some instances significantly alters the subject’s identity, as in the second and third images of the Red Lips edit. Haas et al. [haas2023discovering] are better at preserving the subject’s identity but the edits fail to be localized to the region of interest, changing the subject’s overall expression. Our method succeeds in producing the desired semantic edits in the region of interest with almost imperceptible changes to the other parts of the image.

### v2

@@ 6 Conclusion and Future Work
Appendix A Algorithms & Implementation Details
@@ A.1 Algorithms
Computing Joint and Individual Components in the Latent Space of DMs: Algorithm 1 summarizes our method for decomposing the Jacobian of each region into a joint and an individual component. We start with a real image $\mathbf{x}_{0}$, a denoising network $\mathbf{e}_{\theta}$ and a set of regions $M=\{\mathbf{m}_{1},\cdots,\mathbf{m}_{N}\}$, along with the denoising timestep $t$ and joint and individual ranks $r_{C}$ and $r_{A}$. Starting with $\mathbf{x}_{0}$, we obtain its corresponding noise $\mathbf{x}_{T}$ via DDIM Inversion [Song et al.(2020a)Song, Meng, and Ermon]. We denoise $\mathbf{x}_{T}$ up to timestep $t$ with the standard DDIM reverse process (see Eq. (3) in text). Then we obtain the SVD of the Jacobian of each region with the Subspace Iteration (Algorithm 2). Finally, by utilizing the JIVE (Algorithm 3) we obtain the joint and individual components $\mathbf{C}$ and $\mathbf{A}_{i}$ for each region $i$. The rows of $\mathbf{A}_{i}$ are latent directions that result in meaningful local edits within region $i$.
@@
2:     $\mathbf{x}_{T}\leftarrow\text{DDIMInversion}(\mathbf{x}_{0})$triangleright DDIM Inversion [Song et al.(2020a)Song, Meng, and Ermon]
@@
Jacobian Subspace Iteration: As discussed in Section 4.2 the dimension of the latent space and the output image, result in a Jacobian matrix of approximately 6B parameters. To efficiently calculate the SVD of the Jacobian of the denoising network without storing it in memory we rely on Jacobian Subspace Iteration as proposed by Haas et al. [Haas et al.(2023)Haas, Huberman-Spiegelglas, Mulayoff, and Michaeli]. Note that differently from [Haas et al.(2023)Haas, Huberman-Spiegelglas, Mulayoff, and Michaeli] we calculate the SVD of the Jacobian of the denoising network explicitly within a region of interest $\mathbf{m}$.
@@
Joint and Individual Variation Explained (JIVE) Algorithm: For completeness we provide the iterative JIVE algorithm [Lock et al.(2013)Lock, Hoadley, Marron, and Nobel]. Given a set of $N$ matrices $\mathbf{X}=[\mathbf{X}^{(1)}\cdots\mathbf{X}^{(N)}]$, JIVE iteratively approximates their joint and individual components.
@@ A.2 Impementation Details
Asyrp For Asyrp [Kwon et al.(2022)Kwon, Jeong, and Uh] we use the author’s official code [^4] https://github.com/kwonminki/Asyrp_official. Since Asyrp is a supervised approach we train each attribute with 1k samples following the recipe detailed in their paper. Asyrp uses CLIP to obtain the editing directions and requires a source caption $y^{\text{source}}$ and a target caption $y^{\text{target}}$. The source and target captions we use to obtain the editing directions for each attribute are:
@@
Haas et al. We implement the method of Haas et al. [Haas et al.(2023)Haas, Huberman-Spiegelglas, Mulayoff, and Michaeli] following their paper  [^5] At the time of writing the paper the official codebase of Haas et al. was not public yet.. The editing directions for Smile, Close Eyes, Red Lips, and Gaze edits were obtained from CelebA-HQ samples with index numbers 00009, 00000, 00042, 00003 respectively.
@@
Appendix B Additional Experimental Resuls
@@ B.4 Additional Comparisons
In Figure 10 we present additional qualitative comparisons for two semantic directions not shown in the main paper. Asyrp [Kwon et al.(2022)Kwon, Jeong, and Uh] produces considerable distortions and artifacts in the edited regions and in some instances significantly alters the subject’s identity, as in the second and third images of the Red Lips edit. Haas et al. [Haas et al.(2023)Haas, Huberman-Spiegelglas, Mulayoff, and Michaeli] are better at preserving the subject’s identity but the edits fail to be localized to the region of interest, changing the subject’s overall expression. Our method succeeds in producing the desired semantic edits in the region of interest with almost imperceptible changes to the other parts of the image.